RED company is considering investing $30,000 in machine that is expected to have a useful life of five years and is expected to reduce (save) the firm’s labor costs by $3,850 per year. It is believed that $7,000 can be obtained when the equipment is disposed at the end of year five. Assume that declining balance 200% depreciation method is used.
a. What are the depreciation deduction amounts for years 1 through 5?
b. What is the BV of the asset after 5 years?
Answer #1

initial\ cost,P=\$30,000\\ depreciation,d=\frac{2}{n}=\frac{2}{5}=0.4\\ n=useful\ life\ of\ asset\\

depreciation\ in\ year\ i,D_{i}=P*d*(1-d)^{i-1}\\ book\ value\ in\ year\ i,BV_{i}=P*(1-d)^{i}

Part a)

year depreciation book value
0 30000
1 12000 18000
2 7200 10800
3 4320 6480
4 2592 3888
5 1555.2 2332.8

Part b)

book value of asset afyer 5 years is $ 2332.8